Where will the line 3x+12y=24 cut the y-axis?

Please put the solution.. thank you for being so nice and helpful tutors..^^

Found 2 solutions by Earlsdon, nerdybill:
Answer by Earlsdon(6294)

First get the equation into the "slope-intercept" form: y = mx+b
3x+12y = 24 Subtract 3x from both sides of the equation.
12y = -3x+24 Now divide both sides by 12.
y = (-1/4)x+2 Compare this with the slope-intercept form above:
y = mx+b and you can see that the y-intercept, b = 2 and this is where the line represented by the given equation intercepts (cuts) the y-axis.

Answer by nerdybill(7384)

Where is "cuts the y-axis" is equivalent to the y-intercept.
.
To find it, set x = 0 and solve for y:
3x+12y=24
3(0)+12y=24
0+12y=24
12y=24
y = 2
.
Therefore, it "cuts the y-axis" at (0, 2)